# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ VI.1.4 The non-linear feedback mechanism of the approximation
The self-consistency of the procedure is implicit in eqs.(14-18). By this procedure the nonlinear $g$-dependence of the system Hamiltonian, $H(g)$ is effectively fed back in to the approximating Hamiltonian, $H_{0}$ such that
@@
Thus the above steps generate a mapping from the original ( non-solvable) Hamiltonian $H(g)$ to an exactly solvable AH, $H_{0}$ while preserving the non-linearity of the former in a self-consistent way.
@@ VI.1.5 Features of the Leading order (LO) results in NGAS
The steps which are outlined above are the essential ingredients of the proposed approximation scheme in the leading order(LO). The exact solution of the eigenvalue equation for $H_{0}(g,n)$ given by $E_{0}^{n}(g)$, are naturally identified as the leading-order (LO) result and these are expected to provide the major contribution for the true eigenvalue, $E_{n}(g)$.
At this point, the following observations regarding the approximation scheme are in order:
(i) It is to be noted that in a restricted form, i.e. when the quantum average in eqn.(16) is restricted to the ground state only, the CEQA as expressed in eqn.(16), corresponds to the Hartree-approximation/ mean field approximation in quantum field theory [36] . In this sense NGAS can be regarded as a “generalized” Hartree-approximation method [37]. (ii) It may be emphasized that even the LO results capture the dominant contribution of the (nonlinear) interaction through the self-consistency procedure of NGAS, even though one always deals with an exactly solvable Hamiltonian $H_{0}$. We consider this as a key feature of the approximation method. (iii) Systematic improvement of the LO-result can be implemented in NGAS through the development of an improved perturbaive framework as follows. Consider the following equation which follows from CEQA, eqn.(16):
@@
. Note further that
@@
for arbitrary values of $g$ and $n$, implying that the perturbation contribution remains sub-dominant ( in the sense of QA) irrespective of the coupling strength and excitation levels. It is thus naturally suggested to base a perturbation theory treating $H_{0}$ as the unperturbed Hamiltonian and $H^{\prime}$ as the perturbation. This new perturbation theory is expected to possess features that are remarkably different from those in SFPT. The details are discussed Section-X.
@@ VI.2 Section Summary
In this Section-VI, a non-perturbative general scheme of approximation(NGAS) is formulated which is aimed at retaining the merits but overcoming the difficulties/ limitations of the extant methods/ schemes reviewed in the earlier Section-Ii.
The scheme, NGAS is characterized by the following features/properties:
(a) Universal applicability to a general interacting system in Quantum Theory in the Hamiltonian-formulation irrespective of the strength of interaction.
(b) This scheme, by construction, is shown to be potentially capable of providing accurate approximation for any general interaction of arbitrary strength in quantum theory even in the leading order (LO).
(c) Thus, the scheme is essentially non-perturbative in character in the LO.
(d) Yet, it is indicated in this Section-VI how to improve the LO-results still further through the formulation of an improved perturbation theory in a systematic manner within the scheme while retaining the above features of universality and non-perturbative nature.
(e)The basic input in NGAS is the choice of a suitable approximating-Hamiltonian (AH) ${H_{0}}$ , which is required to be exactly solvable, but which involves certain adjustable (variational) parameters, $\alpha_{i}$. The scheme is implemented by imposing the constraint of equal quantum-average (CEQA) such that the quantum average (QA) of the original Hamiltonian $H$ be equal to that of the AH ${H_{0}}$ with respect to any (arbitrary) eigenstate of the latter. The (variational) minimization of this QA with respect to the undetermined free-parameters completely determines the latter.This is referred as the constraint of Optimality (CO).
(f) Apart from the basic requirements as discussed earlier, there is considerable flexibility in the choice of the AH which may be profitably used to improve accuracy of the results.
(g)The evaluation of the energy eigenvalue of the AH determines the leading order(LO) result in NGAS.
(h) The potential development of an improved perturbation theory for arbitrary strength of interaction is contained in NGAS.
In the following Sections, we implement and illustrate the general approach described above by application to specific bench-mark systems which are traditionally employed as theoretical laboratories to test approximation methods in QT.
@@ VII Application of NGAS in the Leading Order (LO) to Anharmonic Interactions (AHI)
Anharmonic Interactions (AHI) are characterized by potential functions that are polynomial additions to the harmonic-interaction. Examples include: quartic anharmonic and double-well potentials, sextic -, octic- and higher anharmonic potentials ( we confine here to parity-symmetric potentials). These can be represented as:
@@
These interactions are quite basic in QT due to their wide applications (see,e.g.[31]) in diverse areas of Physics, as well as, due to their usefulness in traditionally serving as testing grounds for various approximation schemes(see,e.g.[9]) since exact/accurate numerical results are available(see,e.g [8,9]) for comparison. Besides, extensive literature is accessible for comparison of different methods/schemes of approximation applied to the AHI. These are therefore the primary reasons for application of the NGAS to the AHI.
In the following sections, we describe the application of the NGAS in LO-approximation, to the individual AHI-systems.
@@ VII.1 The Harmonic Approximation
Since the AHI are defined by the modification of the harmonic oscillator potential by addition of further polynomial terms of higher degree, it is natural to choose the approximating Hamiltonian (AH), $H_{0}$ as that for the harmonic oscillator(HO) but possibly with a shift in the field-variable $x$ as well as, in the energy. The frequency, $\omega$ of the HO becomes the other free parameter. The exact expression is provided in the subsections that follow. This choice, then defines the Harmonic Approximation, for the different cases of the AHI.
However, as noted earlier, the NGAS is flexible to the choice of the input AH subject only to the defining properties for the latter. In a later section (Section-VIII), we will be using a different approximation, namely the Infinite- Square-Well-Approximation to illustrate this flexibility.
In the next subsections we consider the different cases of the AHI in turn.
@@ VII.2 The Quartic Anharmonic Oscillator (QAHO)
The quartic-anharmonic oscillator ( QAHO ) is the simplest system  exhibiting  self-interaction. This  system has been extensively investigated leading to a vast amount of literature [8,9]. Its importance arises due to physical applications in areas including field theory, condensed matter physics[83], statistical mechanics [84], non-linear systems [85], classical and quantum chaos [86], inflationary cosmology [87], lattice dynamics, plasma oscillations etc, to cite only a few cases. Besides, the QAHO has also served as a theoretical laboratory  to study convergence of perturbation theory [88], development of non-perturbative approximation methods [89], renormalization [90], vacuum structure [91] and stability analysis [92] etc. This quantum system is the basis in testing various aspects of approximation methods, see e.g. [9-11,17-18,21-23,26-42,47-60].
The Hamiltonian of the system is given by
@@
where ${}g~{}$ is real and positive (We deal with dimensionless variables throughout by using scaling in the standard fashion).
In order to develop the NGAS for the QAHO, we follow the steps outlined in the Section-VI, as described below.
@@
To obtain the exact analytic solution of the spectrum of ${}H_{0}~{}$ we proceed as follows. We choose the “harmonic approximation” for the input approximating Hamiltonian (AH) defined by the following equation
@@
It may be atonce recognized that the AH given by eqn.(25) corresponds to a “shifted” effective harmonic oscillator where the energy is displaced by an amount: $h_{0}$. Note further that the parameter $\omega$ is restricted by physical requirement, to satisfy ${}\omega>~{}0$.
For the implementation of CEQA, one needs to evalute:
$\langle H\rangle=\frac{1}{2}\langle p^{2}\rangle+\frac{1}{2}\langle x^{2}\rangle+g\langle x^{4}\rangle$ (26)
where now the notation is $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle$, for an operator A . Using the definition of $|\phi_{n}\rangle$ as the eigenfunction of $H_{0}$, the relevant operator-averages occurring in $\langle H\rangle$ can be calculated [8,17] by standard methods , e.g. by using the formalism of the ladder operators for $H_{0}$. These are given by the following equations:
@@
$~{}\langle p^{2}\rangle=\omega\xi,$ (27b)
@@
where ${\xi=(n+1/2)}$. Substituting the QA values from eqs.(27a-d) , the QA of the original Hamiltonian $H$ defined by eqn.(26) can now be evaluated as
@@
Carrying out the explicit variational minimization of $\langle H_{0}\rangle$ w.r.t ${}^{\prime}\omega~{}^{\prime}$ i.e $\dfrac{\partial\langle H_{0}\rangle}{\partial\omega}=0$ one obtains the following eqn :
@@
from where $\omega$ can be obtained as the real, +ve root. It may be emphasized at this point that this equation has been derived by several authors [93], but starting from widely different considerations. Similarly, by noting that $h_{0}=E_{0}(g,\omega,\xi)-\omega\xi$ , the expression for $h_{0}$ is obtained and given by
@@
where $f(\xi)\equiv\xi+\frac{1}{4\xi}$ and $\xi=(n+\frac{1}{2})$ ; with the spctral index $n$ taking values : $n=0,1,2,3,.....$. We refer eqn.(29) as the “ gap equation (GE)” and eqn.(30) as the “energy-shift”. As expected, these parameters thus acquire the functional dependence on $g$ and $\xi$:
${}~{}~{}~{}~{}~{}\omega\rightarrow\omega(g,\xi)~{}~{}~{}~{}~{}~{}h_{0}\rightarrow h_{0}(g,\xi)$ (31)
(iii) Determination of the energy-spectrum in the leading order(LO)
Solution of the gap-equation (GE) eqn.(29) constitutes the key ingredient in the calculation of the energy spectrum. From eqs.(28) and (31), it then follows that the eigen values of $H_{0}$ also acquire the required ${}g~{},~{}\xi~{}$-dependence and the leading order (LO) result for the energy is given by
@@
where ‘$\omega$’ is obtained as a solution of the $GE$ for the QAHO given by eqn.(29). It is to be noted that the $GE$, eqn.(29) is in the form of a cubic equation of the type
@@
Then comparing the coefficients of eqs.(33) and (29) the required solution is obtained explicitly as
@@
where, ${}\rho^{-1}~{}=~{}243g^{2}f^{2}(\xi)~{}$. It may be noted that, the solution for ‘${}\omega~{}$’ as given above has the correct limiting behaviour, ${}\omega~{}\longrightarrow~{}1$ for ${}g\longrightarrow 0~{}$ and further that it exhibits the non - analytic dependence on the coupling ’${}g~{}$’ at the orgin characteristic of the non-perturbative nature of the NGAS.
At this point, several features/aspects regarding the leading order (LO) results based on Eqs. (29) - (35) may be noted as follows:
(a) As has been mentioned earlier, equation (29) has been derived independently by several authors [93] starting from widely different considerations.
(b) Moreover, the rigorously established [11] analytic structure of $E(g)$ in the $g$ plane, as well as, the non - analytic dependence on the coupling strength ’ $g$’ , are trivially contained [8,82] already at the LO-level through the solution of eqn. (29), which is explicitly given by eqn.(35).
(c) The dependence of $\omega$ and $h_{0}$ on $\xi$ and $g$ implies that the AH, $H_{0}$ also depends on the same parameters. The physical significance of such dependence is that the eigenfunctions of $H_{0}$, $|\phi_{n}\rangle$ are not mutually orthogonal: $\langle\phi_{m}|\phi_{n}\rangle\neq 0$ for $m\neq n$. (As discussed later, this situation does not affect the development of the perturbation corrections in the standard formulation of perturbation theory (SFPT), which can be generated by a ”wave-function- independent method” , such as the Hyper Virial Theorem (HVT) and the Feynman-Hellman Theorem (FVT)-based formalism employed here ( see, Section-X) .
(d) A further significance of $\omega$ arises from the altered ground state structure [8,50,52] due to interaction, differing non-trivially from the free field ground state. Also, the“dressed ground state” of the interacting system in the LO approximation has lower energy [8,50,52,82] as compared to the “trivial” ground state of the noninteracting-theory for any non-vanishing value of the coupling ’$g$’. This result thus establishes the instability of the free-field ground state in presence of interaction. This aspect has been discussed in detail elsewhere [8,50,52,82].
(e) The $(g,\xi)$-dependence of $\omega$ and $h_{0}$ as determined in the LO, does not get altered later in computing the perturbation corrections at higher orders. We elaborate on this important feature later (Section-X).
(f) Further, as stated earlier, the accuracy of the energy spectrum obtained in the LO, eqn.(3.12) is quite significant- the deviations from the “exact” -results being no more than a few percent [8,50,52,82] over the full-range of $g$ and $n$ This result ensures that the dominant contribution indeed comes from the LO as required in a perturbative-framework. We also display the improvement of results obtained by inclusion of the higher-order corrections to the LO-results, which is discussed in Section-X .
We next apply the method to the case of the quartic-double well oscillator in the following subsection.

### v3

@@ VI.1.4 The non-linear feedback mechanism of the approximation
The concept of self-consistency is embedded within the framework of eqs.(14-18). This approach effectively incorporates the nonlinear $g$-dependence of the system Hamiltonian, $H(g)$, into the approximating Hamiltonian, $H_{0}$, such that
@@
As a result, this process establishes a mapping from the original (non-solvable) Hamiltonian $H(g)$ to an exactly solvable approximating Hamiltonian, $H_{0}$, while preserving the non-linearity of the original Hamiltonian in a self-consistent manner.
@@ VI.1.5 Features of the Leading order (LO) results in NGAS
The steps outlined above form the fundamental ingredients of the proposed approximation scheme at the leading order (LO). The exact solution of the eigenvalue equation for $H_{0}(g,n)$, represented by $E_{0}^{n}(g)$, is naturally recognized as the leading-order (LO) result and is expected to significantly contribute to the true eigenvalue, $E_{n}(g)$. At this stage, the following observations regarding the approximation scheme are noteworthy:   (i) It should be noted that in a restricted form, specifically when the quantum average in eqn.(16) is confined to the ground state only, the CEQA as expressed in eqn.(16) corresponds to the Hartree-approximation or mean field approximation in quantum field theory [36]. In this context, NGAS can be regarded as a ”generalized” Hartree-approximation method [37].
(ii) It is important to emphasize that even the LO results capture the dominant contribution of the (nonlinear) interaction through the self-consistency procedure of NGAS, despite the fact that one is always dealing with an exactly solvable Hamiltonian $H_{0}$. We regard this as a key feature of the approximation method.
(iii) Systematic improvement of the LO-result can be achieved in NGAS through the development of an enhanced perturbative framework as follows. Consider the following equation derived from CEQA, eqn.(16):
@@
It is further noted that
@@
for arbitrary values of $g$ and $n$, indicating that the above correction -  contribution remains sub-dominant (in the sense of QA) regardless of the coupling strength and excitation levels. Consequently, it is naturally suggested to develop a perturbation theory that treats $H_{0}$ as the unperturbed Hamiltonian and $H^{\prime}$ as the perturbation. This novel perturbation theory is anticipated to exhibit characteristics that are significantly distinct from those in SFPT. The details are elaborated in Section-X.
@@ VI.2 Section Summary
In this Section-VI, a novel general scheme of approximation (NGAS) is developed, which seeks to preserve the merits while addressing to overcome the challenges and limitations of the existing methods and schemes discussed in the preceding Section-II.The NGAS is characterized by the following features: (a) Universal applicability to a general interacting system in Quantum Theory within the Hamiltonian formulation, regardless of the interaction strength. (b) By design, this scheme is shown to be potentially capable of providing accurate approximation for any general interaction of arbitrary strength in quantum theory, even at the leading order (LO). (c) Consequently, the scheme is essentially non-perturbative in nature at the LO. (d) Furthermore, this Section-VI indicates how to enhance the accuracy of the LO results through the formulation of an improved perturbation theory in a systematic manner within the scheme, while maintaining the aforementioned features of universality and non-perturbative character. (e) The fundamental input in NGAS is the selection of a suitable approximating-Hamiltonian (AH) H{H}, which must be exactly solvable and involves certain adjustable (variational) parameters, $\alpha_{i}$. The scheme is implemented by imposing the constraint of equal quantum-average (CEQA), ensuring that the quantum average (QA) of the original Hamiltonian $H$ is equal to that of the AH H{H} with respect to any (arbitrary) eigenstate of the latter. The variational minimization of this QA with respect to the undetermined free parameters completely specifies the latter. This is referred to as the constraint of Optimality (CO). (f) Consistent with the basic requirements as discussed earlier, there exists considerable flexibility in the choice of the AH which may be profitably used to improve the overall accuracy of the results . (g)The evaluation of the energy eigenvalue of the AH determines the leading order(LO) result in NGAS. (h) The potential development of an improved perturbation theory for arbitrary strength of interaction is contained in NGAS.
In the following Sections, we implement and illustrate the general approach described above by applying it to specific benchmark systems traditionally employed as theoretical laboratories to test approximation methods in quantum theory (QT).
@@ VII Application of NGAS in the Leading Order (LO) to Anharmonic Interactions (AHI)
Anharmonic Interactions (AHI) are characterized by potential functions that are polynomial additions to the harmonic interaction. Examples include quartic anharmonic and double-well potentials, as well as sextic, octic, and higher-order anharmonic potentials (we confine our discussion here to parity-symmetric potentials). These can be represented as:
@@
These interactions are fundamental in quantum theory due to their extensive applications (see, e.g., [31]) across various domains of physics, as well as their utility in serving as traditional testing grounds for various approximation schemes (see, e.g., [9]), given that exact or highly accurate numerical results are available (see, e.g., [8,9]) for comparison. Furthermore, a substantial body of literature is available for the comparison of different methods and schemes of approximation applied to the AHI. These are the primary reasons for the application of the NGAS to the AHI. In the following sections, we describe the application of the NGAS in LO approximation to the individual AHI systems.
@@ VII.1 The Harmonic Approximation
Given that the AHI are characterized by the modification of the harmonic oscillator potential through the incorporation of additional polynomial terms of higher degree, it is reasonable to select the approximating Hamiltonian (AH), $H_{0}$, as that of the harmonic oscillator (HO), albeit with a potential shift in the field variable $x$ as well as in the energy. The frequency, $\omega$, of the HO becomes the other free parameter. The exact expression is provided in the subsections that follow. This choice then defines the Harmonic Approximation for the different cases of the AHI.
However, as noted earlier, the NGAS is adaptable to the choice of the input AH, subject only to the defining properties of the latter. In a later section (Section-VIII), we will employ a different approximation, namely the Infinite-Square-Well Approximation, to illustrate this flexibility.
In the subsequent subsections, we consider the different cases of the AHI in turn.
@@ VII.2 The Quartic Anharmonic Oscillator (QAHO)
The quartic anharmonic oscillator (QAHO) represents the most fundamental system exhibiting self-interaction. This system has been the subject of extensive research, resulting in a substantial body of literature [8,9]. Its significance is underscored by its applications in various physical domains, including quantum field theory, condensed matter physics [83], statistical mechanics [84], non-linear systems [85], classical and quantum chaos [86], inflationary cosmology [87], lattice dynamics, and plasma oscillations, among others. Furthermore, the QAHO has functioned as a theoretical laboratory for examining the convergence of perturbation theory [88], the development of non-perturbative approximation methods [89], renormalization [90], vacuum structure [91], and stability analysis [92]. This quantum system serves as a foundational model for testing various aspects of approximation methods, as evidenced in numerous studies [9-11,17-18,21-23,26-42,47-60]. The Hamiltonian of the system is given by
@@
where $g~$ is a real and positive parameter (We employ dimensionless variables throughout by utilizing standard scaling methods). To develop the NGAS for the QAHO, we adhere to the steps outlined in Section-VI, as detailed below.
@@
To derive the exact analytical solution of the spectrum of $H_{0}~$, we proceed as follows. We adopt the ”harmonic approximation” for the input approximating Hamiltonian (AH), defined by the equation
@@
It is immediately apparent that the AH given by eqn.(25) corresponds to a ”shifted” effective harmonic oscillator, where the energy is displaced by an amount: $h_{0}$. Furthermore, the parameter $\omega$ is constrained by physical requirements to satisfy $\omega>~0$. For the implementation of CEQA, it is necessary to evaluate:
[…]
where the notation $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle$ is used for an operator A. Utilizing the definition of $|\phi_{n}\rangle$ as the eigenfunction of $H_{0}$, the relevant operator averages in $\langle H\rangle$ can be calculated [8,17] using standard methods, such as the formalism of ladder operators for $H_{0}$. These are given by the following equations:
@@
$\langle p^{2}\rangle=\omega\xi,$ (27b)
@@
where $\xi=(n+1/2)$}. By substituting the QA values from eqs.(27a-d), the QA of the original Hamiltonian $H$ defined by eqn.(26) can now be evaluated as
@@
Conducting the explicit variational minimization of $\langle H_{0}\rangle$ with respect to $^{\prime}\omega~^{\prime}$ i.e., $\dfrac{\partial\langle H_{0}\rangle}{\partial\omega}=0$, yields the following equation:
@@
from which $\omega$ can be determined as the real, positive root. It is noteworthy that this equation has been derived by several authors [93], albeit from significantly different considerations. Similarly, by noting that $h_{0}=E_{0}(g,\omega,\xi)-\omega\xi$, the expression for $h_{0}$ is obtained and given by
@@
where $f(\xi)\equiv\xi+\frac{1}{4\xi}$ and $\xi=(n+\frac{1}{2})$; with the spectral index $n$ taking values: $n=0,1,2,3,..$.... We refer to eqn.(29) as the “gap equation (GE)” and eqn.(30) as the “energy-shift”. As anticipated, these parameters thus acquire a functional dependence on $g$ and $\xi$:
$\omega\rightarrow\omega(g,\xi)~~~~~~h_{0}\rightarrow h_{0}(g,\xi)$ (31)
(iii) Determination of the Energy-Spectrum in the Leading Order (LO) The solution of the gap equation (GE) eqn.(29) constitutes the key component in the calculation of the energy spectrum. From eqs.(28) and (31), it follows that the eigenvalues of $H_{0}$ also acquire the requisite $g~,~\xi~$-dependence, and the leading order (LO) result for the energy is given by
@@
where $\omega$’ is obtained as a solution of the $GE$ for the QAHO given by eqn.(29). It should be noted that the $GE$, eqn.(29), is in the form of a cubic equation of the type
@@
Then, by comparing the coefficients of eqs.(33) and (29), the required solution is obtained explicitly as
@@
It should be noted that the solution for $\omega~$’ as given above exhibits the correct limiting behavior, $\omega~\longrightarrow~1$ for $g\longrightarrow 0~$, and further, it demonstrates the non-analytic dependence on the coupling ’$g~$’ at the origin, characteristic of the non-perturbative nature of the NGAS.
At this juncture, several features pertaining to the leading order (LO) results based on eqs. (29) - (35) warrant attention: (a) As previously mentioned, equation (29) has been independently derived by several authors [93] from diverse independent considerations. (b) Furthermore, the rigorously established [11] analytic structure of $E(g)$ in the $g$-plane, along with the non-analytic dependence on the coupling strength ’$g$’, is inherently present [8,82] at the LO level through the solution of Eq. (29), as explicitly provided by Eq. (35). (c)The dependence of $\omega$ and $h_{0}$ on $\xi$ and $g$ implies that the AH, $H_{0}$, also depends on these parameters. The physical implication of this dependence is that the eigenfunctions of $H_{0}$, $|\phi_{n}\rangle$, are not mutually orthogonal: $\langle\phi_{m}|\phi_{n}\rangle\neq 0$ for $m\neq n$. As discussed later, this situation does not affect the development of perturbation corrections in the standard formulation of perturbation theory (SFPT), which can be generated by a ”wave-function-independent methods,” such as the Hyper Virial Theorem (HVT) and the Feynman-Hellman Theorem (FVT)-based formalism employed here (see, Section-X). (d) An additional significance of $\omega$ arises from the altered ground state structure [8,50,52] due to interaction, which differs non-trivially from the free field ground state. Moreover, the ”dressed ground state” of the interacting system in the LO approximation possesses lower energy [8,50,52,82] compared to the ”trivial” ground state of the noninteracting theory for any non-zero value of the coupling ’$g$’. This result thus establishes the instability of the‘free-field ground state’ in the presence of interaction. This aspect has been discussed in detail elsewhere [8,50,52,82]. (e) The $(g,\xi)$-dependence of $\omega$ and $h_{0}$ as determined in the LO does not undergo alteration in the computation of perturbation corrections at higher orders. We elaborate on this important feature later (Section-X).  (f) Furthermore, as previously stated, the accuracy of the energy spectrum obtained in the LO, Eq. (3.12), is quite significant, with deviations from the ”exact” results being no more than a few percent [8,50,52,82] over the full range of $g$ and $n$. This result ensures that the dominant contribution indeed arises from the LO, as required in a perturbative framework. We also demonstrate the improvement of results obtained by the inclusion of higher-order corrections to the LO results, which is discussed in Section-X. We next apply the method to the case of the quartic-double well oscillator in the following subsection.
Up to this point on Aug 10-2025 (FN)